# The Jacobi Factoring Circuit: Quantum Factoring with Near-Linear Gates and Sublinear Space and Depth

### v1

@@ Abstract
We present a compact quantum circuit for factoring a large class of integers, including some whose classical hardness is expected to be equivalent to RSA (but not including RSA integers themselves). To our knowledge, it is the first polynomial-time circuit to achieve sublinear qubit count for a classically-hard factoring problem; the circuit also achieves sublinear depth and nearly linear gate count. We build on the quantum algorithm for squarefree decomposition discovered by Li, Peng, Du and Suter (Nature Scientific Reports 2012), which relies on computing the Jacobi symbol in quantum superposition. Our circuit completely factors any number $N$, whose prime decomposition has distinct exponents, and finds at least one non-trivial factor if not all exponents are the same. In particular, to factor an $n$-bit integer $N=P^{2}Q$ (with $P$ and $Q$ prime, and $Q<2^{m}$ for some $m$), our circuit uses $\widetilde{O}(m)$ qubits and has depth at most $\widetilde{O}(m+n/m)$, with $\widetilde{O}(n)$ quantum gates. When $m=\Theta(n^{a})$ with $12/3<a<$ 1, the space and depth are sublinear in $n$, yet no known classical algorithms exploit the relatively small size of $Q$ to run faster than general-purpose factoring algorithms. We thus believe that factoring such numbers has potential to be the most concretely efficient classically-verifiable proof of quantumness currently known.
The technical core of our contribution is a new space-efficient quantum algorithm to compute the Jacobi symbol of $A$ mod $B$, in the regime where $B$ is classical and much larger than $A$. Crucially, our circuit reads the bits of the classical value $B$ in a streaming fashion, never storing more than $\widetilde{O}(\log A)$ qubits of quantum information at one time. In the context of the larger Jacobi algorithm for factoring $N=P^{2}Q$, this reduces the overall qubit count to be roughly proportional to the length of $Q$, rather than the length of $N$. Our circuit for computing the Jacobi symbol is also highly gate-efficient and parallelizable, achieving gate count $\widetilde{O}(\log B)$ and depth at most $\widetilde{O}(\log A+\log B/\log A)$. Finally, we note that our circuit for computing the Jacobi symbol generalizes to related problems, such as computing the greatest common divisor, and thus could be of independent interest.
The problem of distinguishing prime numbers from composite numbers and of resolving the latter into their prime factors is known to be one of the most important and useful in arithmetic. [C. F. Gauss, Disquisitiones Arithmeticae, 1801]
@@ 1 Introduction
The difficulty is many-fold; a primary issue is that that the quantum circuits for factoring are still rather large, in terms of gate count, space complexity, and gate depth. For example, Shor’s algorithm seems to require quantum circuits of size $\tilde{O}(n^{2})$ to factor $n$-bit numbers [Sho97] (where the notation $\tilde{O}$ hides factors poly-logarithmic in $n$); a recent improvement by Regev has reduced the asymptotic gate count to $\tilde{O}(n^{3/2})$ [Reg23], but the practical cost of achieving this improved scaling seems rather large [EG24b]. Much effort has been applied to reducing the space-complexity of these circuits, for both Shor [BCDP96, VBE96, Sei01, Cop02, CW00, Bea03, TK06, Zal06, EH17, Gid17, HRS17, Gid19, GE21, KMY24] and Regev’s algorithms [RV24, EG24a], achieving space as low as $\widetilde{O}(n)$ or even $O(n)$ qubits. Yet there seems to be no fundamental obstacle preventing these costs from being improved further, and indeed if we are to have any hope of factoring classically-intractable integers on quantum computers in the near- or medium-term, it will be necessary to do so. Thus we arrive at two questions that are the focus of this paper:
@@
The space and depth complexity are sublinear when $m=\Theta(n^{a})$ with $a<1$. Algorithms to factor numbers of this form have been extensively studied in the cryptography and computational number theory literature [PO96, BDH99, CJLN09, CL09, May10, CFRZ16, HH22, Mul24]; motivated in part by the fact that this problem’s hardness has been used as the basis of several cryptosystems [Oka90, OU98, PT00, Tak98, SS06]. Roughly speaking, to classically factor $N=P^{2}Q$ where $Q$ is the smaller of the two numbers, we have two choices. Either employ a class of special-purpose factoring algorithms along the lines of Lenstra’s elliptic curve method [Len87], which run in time $\exp(\tilde{O}(\sqrt{\log Q}))$; or use the fastest general-purpose factoring algorithm, namely the number field sieve [Pol93, LLMP90, BLP93], which runs in time $\exp(\tilde{O}((\log N)^{1/3}))$. Which one is faster depends on how small $Q$ is relative to $N$. As long as $\log Q=\widetilde{\Omega}((\log N)^{2/3})$, there are no known classical algorithms that exploit the special structure in $N$ to factor faster than general-purpose factoring algorithms. We refer the reader to Section 2.3 for more in-depth discussion on special-purpose factoring.
@@
The Jacobi factoring circuit builds on the remarkable, but apparently little known, work of Li, Peng, Du and Suter [LPDS12] who constructed a quantum circuit to compute the squarefree decomposition of an integer. That is, given as input a positive integer $N$, find the unique $A$ and $B$ such that $N=A^{2}B$ and $B$ is not divisible by the square of any integer (greater than $1$). The quantum part of the LPDS circuit computes a Jacobi symbol mod $N$, followed by a quantum Fourier transform mod $N$. Using the algorithm of Hales and Hallgren [HH00], the quantum Fourier transform mod $N$ can be computed (approximately) with near-linear size quantum circuits. We observe that using Schönhage’s GCD algorithm [Sch71, BS96, Möl08], Jacobi symbols can be computed in near-linear time as well. Overall, this gives a near-linear size (and also near-linear space and depth) quantum circuit for squarefree decomposition which, in particular, factors numbers $N=P^{2}Q$ where $P,Q$ are prime numbers.
@@
  • Our final contribution, presented in Section 5, is the observation that an algorithm for squarefree decomposition suffices to completely factor a general class of integers of inverse-polynomial density, namely any integer whose prime factorization has distinct exponents.
@@
Our result can be seen to complement the classical factoring algorithms, e.g. [Len87, Mul24] that exploit various types of structure. The result by [LPDS12] takes a first step in this direction by showing that some integers with special structure (e.g. $N=P^{2}Q$ with $P,Q$ prime) become easier to factor quantumly. Our result builds on this, demonstrating that these integers can be quantumly factored even more easily (i.e. in much lower space and depth) if $\log Q\ll\log N$. We emphasize that, as long as $\log Q\geq\widetilde{\Omega}((\log N)^{2/3})$, this structure cannot be exploited by any known special-purpose classical factoring algorithms (we discuss this more in Section 2.3).
In contrast, prior quantum factoring algorithms [Sho97, Reg23] do not seem to benefit from any such structure. With that said, an important open question is whether our algorithm can be leveraged or extended to factor integers in general (we note that it would suffice to devise a way to factor squarefree integers; see Sections 3 and 5 for details).
@@
  • Factoring algorithms: Shor’s algorithm for factoring [Sho97], when implemented with a low-depth quantum multiplication circuit [NZLS23], costs $\widetilde{O}(n^{2})$ gates, $\widetilde{O}(n)$ qubits, and $\widetilde{O}(n)$ depth. [^1] There also exist log-depth implementations of Shor’s algorithm [CW00], but they come at the cost of far worse gate and qubit counts. We include asymptotics for this circuit in Table 1 for completeness. Regev’s recent improved factoring algorithm [Reg23], together with the optimizations of [RV24] (and using the same low-depth multiplier), can be implemented in $\widetilde{O}(n^{1.5})$ gates, $\widetilde{O}(n)$ qubits, and $\widetilde{O}(n^{0.5})$ depth. The previously proposed Jacobi factoring circuit [LPDS12], together with the algorithm by [Sch71] for computing Jacobi symbols, uses $\widetilde{O}(n)$ gates, space, and depth.
@@ 1.1 The LPDS Circuit for Squarefree Decomposition
With quantum period finding in mind, this naturally suggests the following procedure: (1) start with a uniform superposition over all $x$ mod $N$; (2) compute and measure the Jacobi symbol $\left(\frac{x}{N}\right)$; and (3) use Shor’s period-finding procedure [Sho97] to recover $Q$. The apparent obstacle is that once we measure the Jacobi symbol $\left(\frac{x}{N}\right)$, we will end up not with one periodic signal modulo $Q$ but a superposition of several periodic signals modulo $Q$.
@@
$\frac{1}{\sqrt{N}}\cdot\sum_{x\in[0,N-1]}\ket{x}\stackrel{{\scriptstyle\mbox{\small measure $\mathsf{manyJac}_{N,k}$}}}{{\relbar\joinrel\relbar\joinrel\relbar\joinrel\relbar\joinrel\relbar\joinrel\relbar\joinrel\relbar\joinrel\relbar\joinrel\relbar\joinrel\relbar\joinrel\relbar\joinrel\relbar\joinrel\relbar\joinrel\relbar\joinrel\rightarrow}}\frac{1}{P}\cdot\sum_{j\in[0,P^{2}-1]}\ket{x_{0}+jQ}\stackrel{{\scriptstyle\mbox{\small QFT mod $N$}}}{{\longrightarrow}}\frac{1}{\sqrt{Q}}\sum_{j\in[0\ldots Q-1]}e^{-2\pi ix_{0}j/Q}\bigg{\lvert}\,\frac{jN}{Q}\bigg{\rangle}$ (1)
@@
This would, however, result in a rather large circuit: to uniquely fix $x$ mod $Q$, one would certainly need to compute at least $\Omega(\log Q)$ Jacobi symbols (and perhaps even a larger $\mathsf{poly}(\log Q)$ [BL96, CW24]). The key result of [LPDS12] is that just computing and measuring a single Jacobi symbol already suffices (even though we would be working with a superposition of periodic signals333[HH00] provides a black-box algorithm for finding the period of “many-to-one” periodic functions like this, however it requires a super-constant number of calls to the Fourier sampling subroutine. The Gauss sum analysis of [LPDS12] (and that of the present work) provides much better efficiency, showing that just one iteration of Fourier sampling suffices to find the period with probability $\Omega(1)$.), shown using a remarkable analysis involving Gauss sum bounds! Indeed, the QFT of the signal after measuring a single Jacobi symbol, namely that of $x$ mod $N$, will be very similar to the end result in equation 1 except that each basis state will receive a sum of several amplitudes. In particular, if we measure the Jacobi symbol $\left(\frac{x}{N}\right)$ and obtain a value $b\in\left\{-1,1\right\}$, [^4] We once again assume here that we do not obtain the relatively unlikely outcome of $b=0$; we will handle this more carefully in the relevant technical sections. each non-zero basis state |$\ket{jN/Q}$ will have absolute amplitude
$\approx\frac{1}{Q}\cdot\left|\sum_{x_{0}:\left(\frac{x_{0}}{Q}\right)=b}\exp\left(-\frac{2\pi ix_{0}j}{Q}\right)\right|$
By standard Gauss sum bounds (see Section 2.4.2 for details), the summation is lower-bounded by $\Omega(\sqrt{Q})$, and so we know that each non-zero basis state will have amplitude $\Omega(1/\sqrt{Q})$. Since there are $Q-1$ such states, a measurement will give us a non-zero multiple of $N/Q=P^{2}$ with a constant probability (in fact, [LPDS12] shows that this probability can be made to equal 1!).
@@ 1.2 Technical Overview
The reason this does not immediately suffice for our setting is that we are not working with one periodic signal; we would be working with a superposition of periodic signals corresponding to values $x_{0}\in[0,Q-1]$ that have Jacobi symbol $\left(\frac{x_{0}}{Q}\right)=b$ for some $b\in\left\{-1,1\right\}$. To get around this, we combine elements of Shor’s analysis [Sho97] with the Gauss sum analysis introduced by [LPDS12].
@@
The computational bottleneck in the [LPDS12] factoring circuit is computing the Jacobi symbol $\left(\frac{x}{N}\right)$, where $x\in[0,N-1]$ is in superposition. This can be done in gates (and hence space/depth) $\widetilde{O}(n)$ [Sch71, BS96, Möl08], where $n$ is the number of bits in $N$, which is a near-linear gate complexity and hence essentially tight.
However, thanks to our first contribution, we need only compute $\left(\frac{x}{N}\right)$ for $x\leq\mathsf{poly}(Q)$; moreover, since $N$ is classically known, there is the tantalizing possibility that the number of qubits could be pushed down to linear in $\log Q$ rather than $\log N$. We show that this is indeed the case (up to polylogarithmic factors), by constructing a quantum circuit that computes the Jacobi symbol using space $\widetilde{O}(m)$ qubits for any $m\geq\log Q$. Our circuit is also very efficient, achieving gate count $\widetilde{O}(n)$, parallelized into depth $\widetilde{O}(n/m+m)$. [^5] If $\log Q<O((\log N)^{1/2})$, the depth and space cannot both be made to both scale with $\log Q$ simultaneously, because the space-time volume (space times depth) is lower bounded by the gate count, and the gate count is lower bounded by $O(\log N)$. However, the parameters can be tuned to achieve a continuous tradeoff between the two, while maintaining a space-time product nearly linear in $\log N$.
@@
  • Finally, there is an algorithm due to Schönhage [Sch71, BS96, Möl08] that runs in $\widetilde{O}(\log N)$ gates, but does not come with any better guarantees on the space and depth.
@@
The challenge, and room for creativity, is in implementing step 1. The “extended Euclidean” approach would be to compute $N\bmod{x}$, but it is not clear how to do this reversibly in low space and depth. Instead, we find a multiple $kx$ of $x$ such that both of the following are true: (a) $N-kx$ is divisible by $2^{n-m}$; and (b) $kx<2^{n}$. For now one should consider $m=\lceil\log x\rceil$; in some cases, one may choose $m\geq\lceil\log x\rceil$ to improve efficiency, as we discuss later. To be explicit, this allows us to compute the Jacobi symbol via the following chain of transformations:
@@
Section 5: Fully Factoring Special Integers.
Finally, we present a black-box reduction implying that any algorithm for squarefree decomposition can be used to fully factor integers $N$ with distinct exponents in their prime factorization; i.e. $N$ that can be written as $p_{1}^{\alpha_{1}}\ldots p_{r}^{\alpha_{r}}$ for distinct primes $p_{1},\ldots,p_{r}$ and distinct positive exponents $\alpha_{1},\ldots,\alpha_{r}$. Such integers have been studied before and are referred to as special integers [AM17]. The idea is simple: using the algorithm for squarefree decomposition, we can recover

### v3

@@ Abstract
We present a compact quantum circuit for factoring a large class of integers, including some whose classical hardness is expected to be equivalent to RSA (but not including RSA integers themselves). Most notably, we factor $n$-bit integers of the form $P^{2}Q$ with $\log Q=\Theta(n^{a})$ for $a\in(2/3,1)$ in space and depth sublinear in n (specifically, $\widetilde{O}(\log Q)$) using $\widetilde{O}(n)$ quantum gates; for these integers, no known classical algorithms exploit the relatively small size of $Q$ to run asymptotically faster than general-purpose factoring algorithms. To our knowledge, this is the first polynomial-time circuit to achieve sublinear qubit count for a classically-hard factoring problem. We thus believe that factoring such numbers has potential to be the most concretely efficient classically-verifiable proof of quantumness currently known.
Our circuit builds on the quantum algorithm for squarefree decomposition discovered by Li, Peng, Du, and Suter (Nature Scientific Reports 2012), which relies on computing the Jacobi symbol in quantum superposition. The technical core of our contribution is a new space-efficient quantum algorithm to compute the Jacobi symbol of $A$ mod $B$, in the regime where $B$ is classical and much larger than $A$. Our circuit for computing the Jacobi symbol generalizes to related problems such as computing the greatest common divisor and modular inverses, and thus could be of independent interest.
@@ 1 Introduction
The difficulty is many-fold; a primary issue is that that the quantum circuits for factoring are still rather large, in terms of gate count, space complexity, and gate depth. For example, Shor’s algorithm seems to require quantum circuits of size $\tilde{O}(n^{2})$ to factor $n$-bit numbers [Sho97] (where the notation $\tilde{O}$ hides factors poly-logarithmic in $n$); a recent improvement by Regev has reduced the asymptotic gate count to $\tilde{O}(n^{3/2})$ per run  [Reg25], but the practical cost of achieving this improved scaling seems rather large [EG24b]. Much effort has been applied to reducing the space-complexity of these circuits, for both Shor [BCDP96, VBE96, Sei01, Cop02, CW00, Bea03, TK06, Zal06, EH17, Gid17, HRS17, Gid19, GE21, KMY24] and Regev’s algorithms [RV24, EG24a], achieving space as low as $\widetilde{O}(n)$ or even $O(n)$ qubits. Particularly notable is a recent work which reduced the space cost of Shor’s algorithm to $n/2+o(n)$ qubits [CFS24]. Yet there seems to be no fundamental obstacle preventing these costs from being improved further, and indeed if we are to have any hope of factoring classically-intractable integers on quantum computers in the near- or medium-term, it will be necessary to do so. Thus we arrive at two questions that are the focus of this paper:
@@
The space and depth complexity are sublinear when $m=\Theta(n^{a})$ with $a\in(0,1)$. Algorithms to factor numbers of this form have been extensively studied in the cryptography and computational number theory literature [PO96, BDH99, CJLN09, CL09, May10, CFRZ16, HH22, Mul24]; motivated in part by the fact that this problem’s hardness has been used as the basis of several cryptosystems [Oka90, OU98, PT00, Tak98, SS06]. Roughly speaking, to classically factor $N=P^{2}Q$ where $Q$ is the smaller of the two numbers, we have two choices. Either employ a class of special-purpose factoring algorithms along the lines of Lenstra’s elliptic curve method [Len87], which run in time $\exp(\tilde{O}(\sqrt{\log Q}))$; or use the fastest general-purpose factoring algorithm, namely the number field sieve [Pol93, LLMP90, BLP93], which runs in time $\exp(\tilde{O}((\log N)^{1/3}))$. Which one is faster depends on how small $Q$ is relative to $N$. As long as $\log Q=\widetilde{\Omega}((\log N)^{2/3})$, there are no known classical algorithms that exploit the special structure in $N$ to factor faster than general-purpose factoring algorithms. We refer the reader to Section 2.3 for more in-depth discussion on special-purpose factoring.
@@
The Jacobi factoring circuit builds on the remarkable, but apparently little known, work of Li, Peng, Du and Suter [LPDS12] who constructed a quantum circuit to compute the squarefree decomposition of an integer. That is, given as input a positive integer $N$, find the unique $A$ and $B$ such that $N=A^{2}B$ and $B$ is not divisible by the square of any integer (greater than $1$). The quantum part of the LPDS circuit computes a Jacobi symbol mod $N$, followed by a quantum Fourier transform mod $N$. Using the algorithm of Hales and Hallgren [HH00], the quantum Fourier transform mod $N$ can be computed (approximately) with near-linear size quantum circuits. We observe that using Schönhage’s GCD algorithm [Sch71, TY90, BS96, Möl08], Jacobi symbols can be computed in near-linear time as well. Overall, this gives a near-linear size (and also near-linear space and depth) quantum circuit for squarefree decomposition which, in particular, factors numbers $N=P^{2}Q$ where $P,Q$ are prime numbers.
@@
  • Our final contribution, presented in Section 5, is the observation that an algorithm for squarefree decomposition suffices to completely factor a general class of integers of inverse-polynomial density, namely any integer whose prime factorization has distinct exponents. A similar observation is well-known in the context of factoring polynomials [Yun76], and the high-level ideas are similar between the two settings.
@@
Our result can be seen to complement the classical factoring algorithms, e.g. [Len87, Mul24] that exploit various types of structure. The result by [LPDS12] takes a first step in this direction by showing that some integers with special structure (e.g. $N=P^{2}Q$ with $P,Q$ prime) become polynomially easier to factor quantumly. Our result builds on this, demonstrating that these integers can be quantumly factored even more easily (i.e. in much lower space and depth) if $\log Q\ll\log N$. We emphasize that, as long as $\log Q\geq\widetilde{\Omega}((\log N)^{2/3})$, this structure cannot be exploited by any known special-purpose classical factoring algorithms (we discuss this more in Section 2.3).
In contrast, prior quantum factoring algorithms [Sho97, Reg25] do not seem to benefit polynomially from any such structure. Remarkable algorithms by [EH17, CFS24] that achieve constant-factor improvements over Shor’s original construction depend upon the input being an RSA integer (the product of two primes of equal bit length), but no prior algorithms have better asymptotic scaling depending on the structure of the input. A related, and important, open question is whether our algorithm can be leveraged or extended to factor integers in general (we note that it would suffice to devise a way to factor squarefree integers; see Sections 3 and 5 for details).
Another important future direction is classical cryptanalysis for factoring integers of the form $P^{2}Q$ where $Q$ is much smaller than $P$, since this is the regime in which we get sublinear space. To the best of our knowledge, existing algorithms [Len87, BDH99, Mul24] do not offer any significant improvements in this regime, but this is not a regime that was previously of much practical interest, so we leave further investigation along these lines to future work.
@@
  • Factoring algorithms: Shor’s algorithm for factoring [Sho97], when implemented with a low-depth quantum multiplication circuit [NZLS23], costs $\widetilde{O}(n^{2})$ gates, $\widetilde{O}(n)$ qubits, and $\widetilde{O}(n)$ depth. [^1] There also exist log-depth implementations of Shor’s algorithm [CW00], but they come at the cost of far worse gate and qubit counts. We include asymptotics for this circuit in Table 1 for completeness. Regev’s recent improved factoring algorithm [Reg25], together with the optimizations of [RV24] (and using the same low-depth multiplier), can be implemented in $\widetilde{O}(n^{1.5})$ gates, $\widetilde{O}(n)$ qubits, and $\widetilde{O}(n^{0.5})$ depth. Note, however, that this is the gate count per run, and as $O(n^{0.5})$ runs are required, Regev does not improve total gate count across runs. The previously proposed Jacobi factoring circuit [LPDS12], together with the algorithm by [Sch71] for computing Jacobi symbols, uses $\widetilde{O}(n)$ gates, space, and depth.
@@ 1.1 The LPDS Circuit for Squarefree Decomposition
With quantum period finding in mind, this naturally suggests the following procedure: (1) start with a uniform superposition over all $x$ mod $N$; (2) compute and measure the Jacobi symbol $\left(\frac{x}{N}\right)$; and (3) use Shor’s period-finding procedure [Sho97] to recover $Q$. The apparent obstacle is that once we apply a phase of $\left(\frac{x}{N}\right)$, we will end up not with one periodic signal modulo $Q$ but a superposition of several periodic signals modulo $Q$.
@@
$\frac{1}{\sqrt{N}}\cdot\sum_{x\in[0,N-1]}\ket{x}\stackrel{{\scriptstyle\mbox{\small measure $\mathsf{manyJac}_{N,k}$}}}{{\relbar\joinrel\relbar\joinrel\relbar\joinrel\relbar\joinrel\relbar\joinrel\relbar\joinrel\relbar\joinrel\relbar\joinrel\relbar\joinrel\relbar\joinrel\relbar\joinrel\relbar\joinrel\relbar\joinrel\relbar\joinrel\rightarrow}}\frac{1}{P}\cdot\sum_{j\in[0,P^{2}-1]}\ket{x_{0}+jQ}\stackrel{{\scriptstyle\mbox{\small QFT mod $N$}}}{{\longrightarrow}}\frac{1}{\sqrt{Q}}\sum_{j\in[0,Q-1]}e^{-2\pi ix_{0}j/Q}\bigg{\lvert}\,\frac{jN}{Q}\bigg{\rangle}$ (1)
@@
This would, however, result in a rather large circuit: to uniquely fix $x$ mod $Q$, one would certainly need to compute at least $\Omega(\log Q)$ Jacobi symbols (and perhaps even a larger $\mathsf{poly}(\log Q)$ [BL96, CW24]). The key result of [LPDS12] is that just computing and measuring a single Jacobi symbol already suffices (even though we would be working with a superposition of periodic signals333[HH00] provides a black-box algorithm for finding the period of “many-to-one” periodic functions like this, however it requires a super-constant number of calls to the Fourier sampling subroutine. The Gauss sum analysis of [LPDS12] (and that of the present work) provides much better efficiency, showing that just one iteration of Fourier sampling suffices to find the period with probability $\Omega(1)$.). Even better, instead of measuring the Jacobi symbol we can simply apply a phase equal to the Jacobi symbol $\left(\frac{x}{N}\right)$. The effect of this is to eliminate any amplitude placed by the post-QFT state on |$\ket{0}$ (which is not helpful for factoring $N$). Indeed, the QFT of the signal after measuring a single Jacobi symbol, namely that of $x$ mod $N$, will be very similar to the end result in Equation (1) except that each basis state will receive a sum of several amplitudes. In particular, if we apply a phase equal to the Jacobi symbol $\left(\frac{x}{N}\right)$ [^4] We once again assume here that the Jacobi symbol is always in $\left\{-1,1\right\}$ even though it can occasionally also be 0; we will handle this more carefully in the relevant technical sections. each non-zero basis state |$\ket{jN/Q}$ will have absolute amplitude
$\approx\frac{1}{Q}\cdot\left|\sum_{x_{0}\in[1,Q-1]}\left(\frac{x_{0}}{Q}\right)\exp\left(-\frac{2\pi ix_{0}j}{Q}\right)\right|.$
By standard Gauss sum bounds (see Section 2.4.2 and Remark 3 for details), the summation is lower-bounded by $\Omega(\sqrt{Q})$, and so we know that each non-zero basis state will have amplitude $\Omega(1/\sqrt{Q})$. Since there are $Q-1$ such states, a measurement will give us a non-zero multiple of $N/Q=P^{2}$ with a constant probability (in fact, [LPDS12] shows that this probability is 1!).
@@ 1.2 Technical Overview
The reason this does not immediately suffice for our setting is that we are not working with one periodic signal; we would be working with a superposition of periodic signals corresponding to values $x_{0}\in[0,Q-1]$ grouped according to their Jacobi symbol $\left(\frac{x_{0}}{Q}\right)$. To get around this, we combine elements of Shor’s analysis [Sho97] with the Gauss sum analysis introduced by [LPDS12].
@@
The computational bottleneck in the [LPDS12] factoring circuit is computing the Jacobi symbol $\left(\frac{x}{N}\right)$, where $x\in[0,N-1]$ is in superposition. This can be done in gates (and hence space/depth) $\widetilde{O}(n)$ [Sch71, TY90, BS96, Möl08], where $n$ is the number of bits in $N$, which is a near-linear gate complexity and hence essentially tight.
However, thanks to our first contribution, we need only compute $\left(\frac{x}{N}\right)$ for $x\leq\mathsf{poly}(Q)$; moreover, since $N$ is classically known, there is the tantalizing possibility that the number of qubits could be pushed down to linear in $\log Q$ rather than $\log N$. We show that this is indeed the case (up to polylogarithmic factors), by constructing a quantum circuit that computes the Jacobi symbol using space $\widetilde{O}(m)$ qubits for any $m\geq\log Q$. Our circuit is also very efficient, achieving gate count $\widetilde{O}(n)$, parallelized into depth $\widetilde{O}(n/m+m)$. [^5] If $\log Q<O((\log N)^{1/2})$, the depth and space cannot both be made to scale with $\log Q$ simultaneously, because the space-time volume (space times depth) is lower bounded by the gate count, and the gate count is lower bounded by $O(\log N)$. However, the parameters can be tuned to achieve a continuous tradeoff between the two, while maintaining a space-time product nearly linear in $\log N$.
@@
  • Finally, there is an algorithm due to Schönhage [Sch71, TY90, BS96, Möl08] that runs in $\widetilde{O}(\log N)$ gates, but does not come with any better guarantees on the space and depth.
@@
The challenge, and room for creativity, is in implementing step 1. To do this, we find a multiple $kx$ of $x$ such that both of the following are true: (a) $N-kx$ is divisible by $2^{n-m}$; and (b) $kx<2^{n}$. For now one should consider $m=\lceil\log x\rceil$; in some cases, one may choose $m\geq\lceil\log x\rceil$ to improve efficiency, as we discuss later. To be explicit, this allows us to compute the Jacobi symbol via the following chain of transformations:
@@
A conceptually simpler but concretely less direct and efficient variant [^6] We thank Daniel J. Bernstein for suggesting this perspective on our algorithm. of our algorithm essentially follows the “extended Euclidean” blueprint: we use $kx$ to compute $N\bmod{x}=\left[\left(\frac{N-kx}{2^{n-m}}\right)\cdot\left(2^{n-m}\bmod{x}\right)\right]\bmod{x}$. Then we can transform $\left(\frac{N}{x}\right)\rightarrow\left(\frac{N\bmod{x}}{x}\right)$ and finish from there using [Sch71]. In practice, this is unnecessary extra computation, so we proceed using our more specialized blueprint and do not bother with computing $N\bmod{x}$. One could of course attempt to directly compute $N\bmod{x}$ as usual using long division, but it is not clear how to do this reversibly in low space and depth.
@@
Our algorithm builds on Montgomery reduction [Mon85] and the binary GCD algorithm and can be thought of as a “reversed” variant of long division; long division starts from the most significant bit (MSB) and iterates towards the least significant bit (LSB) to find a multiple of $x$ that agrees with $N$ in the MSBs, whereas we start from the LSB and iterate towards the MSB to find a multiple of $x$ that agrees with $N$ in the LSBs. The benefit of proceeding in this reversed way is that uncomputing intermediate states now becomes easy, by way of a simple comparison that only depends on the MSBs of our intermediate state.
Section 5: Completely Factoring Special Integers.
Finally, we present a black-box reduction implying that any algorithm for squarefree decomposition can be used to completely factor integers $N$ with distinct exponents in their prime factorization; i.e. $N$ that can be written as $p_{1}^{\alpha_{1}}\ldots p_{r}^{\alpha_{r}}$ for distinct primes $p_{1},\ldots,p_{r}$ and distinct positive exponents $\alpha_{1},\ldots,\alpha_{r}$. Such integers have been studied before and are referred to as special integers [AM17] — in fact, these have even been proposed for use in cryptographic applications [Sch18].
A similar observation is well-known in the context of factoring polynomials; the motivation for this is that if a univariate polynomial $f$ is divisible by the square of some (non-constant) polynomial, this can be detected easily by taking the GCD of $f$ and its derivative $f^{\prime}$. Yun [Yun76] shows that this can be extended to decompose any polynomial $f(x)$ into a factorization $f(x)=g_{1}(x)^{\alpha_{1}}\ldots g_{r}(x)^{\alpha_{r}}$ where the $g_{i}$’s are squarefree and pairwise coprime and the $\alpha_{i}$’s are distinct. Many polynomial factorization algorithms thus begin with this step as a subroutine.
Unsurprisingly, the algorithm in our setting bears some high-level similarity to Yun’s algorithm [Yun76] and is based on a simple idea: using the algorithm for squarefree decomposition, we can recover